40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 10 15:41:37 2019, aaron, Update, IOO, WFS measurements Screenshot_from_2019-09-10_18-51-28.pngmctrend_190910.png
    Reply  Wed Sep 11 10:26:56 2019, aaron, Update, IOO, WFS measurements 
       Reply  Wed Sep 11 14:37:43 2019, aaron, Update, IOO, WFS measurements 20190911_WFS.jpg20190911_WFS_2.jpg
          Reply  Thu Sep 12 12:42:31 2019, aaron, Update, IOO, WFS measurements WFS2_TFs.pdfD755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg
             Reply  Fri Sep 13 10:36:03 2019, aaron, Update, IOO, WFS measurements TF-.pngWFS2_TFs.pdf
                Reply  Mon Sep 16 12:38:59 2019, aaron, Update, IOO, WFS measurements 
                Reply  Thu Oct 3 11:38:35 2019, aaron, Update, IOO, WFS measurements 
          Reply  Mon Sep 30 13:20:55 2019, aaron, Update, IOO, shot noise measurement 
       Reply  Fri Sep 13 10:53:40 2019, aaron, Update, IOO, WFS loop measurements 
          Reply  Mon Sep 16 05:08:04 2019, rana, Update, IOO, WFS loop measurements LSCsensingMatrix.ipynbASCsensingMatrix.ipynb
             Reply  Mon Sep 16 11:55:58 2019, rika, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 09:41:48 2019, gautam, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 10:47:44 2019, rika, Update, IOO, WFS loop measurements MC2.pdf
                   Reply  Wed Sep 18 14:45:52 2019, rika, Update, IOO, WFS loop measurements Screenshot_from_2019-09-18_18-15-34.png
             Reply  Tue Sep 17 10:34:48 2019, aaron, Update, IOO, WFS loop measurements 
                Reply  Tue Oct 8 20:39:42 2019, aaron, Update, IOO, WFS loop measurements 
                   Reply  Wed Oct 9 09:37:28 2019, aaron, Update, IOO, WFS loop measurements 
Message ID: 14868     Entry time: Tue Sep 10 15:41:37 2019     Reply to this: 14871
Author: aaron 
Type: Update 
Category: IOO 
Subject: WFS measurements 

[rika, aaron, rana]

We are getting the MC locked in anticipation of making some WFS transfer function measurements.

The PSL screen was all white boxes, so I keyed the PSL crate and burt restored the settings from 11:19am Sep 5 (somewhat earlier than we started rebooting computers). Following this, I ran Milind's unstick.py and then the PSL autolocker script; both worked on the first go, great work Milind!

The modecleaner autolocking script is having substantially more trouble. Rana found that pitch and yaw sliders for all MC optics have been swapped--we think it's because the camera at MC2 has been rotated. Note that for now, sliding pitch gives a change in yaw, and sliding yaw changes pitch.

Improving MC alignment

We noticed that with the WFS servo on, the modecleaner would be well aligned for a while (MC trans ~ 14000), only to lose lock after several minutes. We held the MC2_TRANS_PIT/YAW outputs at 0, so the MC2 QPD does not affect the WFS loop; the beam is well centered on WFS1/2, but not on the MC2 QPD, and with this signal out of the loop MC TRANS recovers to ~15000 counts (consistent with the quiet times over the last 90 days, see attachment 2). Attachment 1 shows the MC lock degrading, followed by some noise where we lost lock, and finally a visible increase in MC trans when we remove the MC2 QPD from the WFS loop.

mode cleaner alignment setting

MC1 Pich 4.4762     Yow 4.4669

MC2 Pich 3.7652     Yow -1.5482

MC3 Pich -0.4159    Yow 1.1477

 

After automatic locking MC, we stopped automatical locking and took alignment to the center of QPD.

And then again did the automatic locking MC. Finaly Rana move to best alignment.

 

Mode cleaner Alignment Setting

MC1 Pich 4.4942   Yow 4.6956

MC2 Pich 3.7652   Yow -1.5600

MC3 Pich -0.3789   Yow 1.1477

 

Measured sine response

We used diaggui to measure the response of WFS1/WFS2/MC2 pitch (yaw) to excitations in MC1/MC2/MC3 pitch (yaw). Seeing fluctuations of amplitude ~1 on the MCX_PIT/YAW_OUT channels, we used an amplitude 0.01 excitation at 20 Hz. We will work on scripting some of this tomorrow.

 

 

Attachment 1: Screenshot_from_2019-09-10_18-51-28.png  56 kB  Uploaded Tue Sep 10 20:43:10 2019  | Hide | Hide all
Screenshot_from_2019-09-10_18-51-28.png
Attachment 2: mctrend_190910.png  30 kB  Uploaded Tue Sep 10 20:45:04 2019  | Hide | Hide all
mctrend_190910.png
ELOG V3.1.3-